Dear tutors please help me with my bring home exam..,!! pleasE!!
1. Find all sets of three consecutive integers whose sum is greater that 54 and less than 63.

Answer by checkley71(8403)

55=x+x+1+x+2 or 3x=53 or x=53/3 or 17 2/3 thus 18+19+20=57 &
19+20+21=60 & 20+21+22=63 reject because these three =63
thus there are two sets that meet the requirements & they are (18,19,20) & (19,20,21)